OpenStudy (anonymous):

log3 7+log3 6-log3 2

OpenStudy (john_es):

Do you mean, \[\log_37+\log_36-\log_32\]?

OpenStudy (john_es):

Do you remember the properties, \[\log a+\log b=\log(a\cdot b)\]\[\log a-\log b=\log(a/ b)\]\[\log_a (a^n)=n\]